OpenStudy (anonymous):

A pendulum bob swings 4.0 cm on its first oscillation. On each subsequent oscillation the bob travels 3/4 the previous distance. Find the total distance the bob travels before coming to a rest.

OpenStudy (dumbcow):

geometric series... A_n = 4*(3/4)^(n-1) find infinite sum of series --> S_n = 4/(1-3/4)

OpenStudy (anonymous):

\[a+ar+ar^2+ar^3+...=\frac{a}{1-r}\] so long as \(-1<r<1\) in your case \(a=4,r=\frac{3}{4}\)

OpenStudy (anonymous):

so -.5 or -1/2?

OpenStudy (anonymous):

neither the problem asks "what is the total distance" so the answer must be positive did you calculate \(\frac{4}{1-\frac{3}{4}}\) ?

OpenStudy (anonymous):

oh, sorry i was going off the first guys equation. so its -.16?

OpenStudy (anonymous):

hold on lets go slow

OpenStudy (anonymous):

16, not .16 sorry

OpenStudy (anonymous):

the answer is clearly a positive number right? it is asking for a distance "Find the total distance ..."

OpenStudy (anonymous):

ah yes, it is 16

OpenStudy (anonymous):

great thanks

OpenStudy (anonymous):

yw